# MATHWELL: Generating Educational Math Word Problems at Scale

### v3

@@ 4 Experiments
We evaluate MATHWELL using both human evaluations with domain experts and automatic evaluations. We compare 250 randomly-selected MATHWELL outputs to those of LLEMMA (34B) Azerbayev et al. (2023), MAmmoTH (70B) Yue et al. (2023), and Llama-2 (70B) Touvron et al. (2023), which are recent and popular. LLEMMA and MAmmoTH represent two different approaches to training math LLMs, as MAmmoTH is trained as a QA model Yue et al. (2023), while LLEMMA is trained more generally using a corpus of math content Azerbayev et al. (2023). Llama-2 serves as a baseline to ensure task-specific finetuning improves performance. Each model is prompted using examples from SGSM and asked to create a question/answer pair on a random topic K-8 students are interested in. Our 250 evaluation sample for each model is roughly 2.5-5 times larger than the human evaluation samples used in other word problem generator studies Jiao et al. (2023); Koncel-Kedziorski et al. (2016); Niyarepola et al. (2022); Qin et al. (2024); Zhou and Huang (2019); Zhou et al. (2023); Zong and Krishnamachari (2023).
@@ 4.1 Human Evaluation
Solvable Questions | MaC Questions | |||||||||||||
Model | Add. | Sub. | Mult. | Div. | Frac. | Dec. | No Ops | Add. | Sub. | Mult. | Div. | Frac. | Dec. | Total Ops |
LLEMMA | 34.4 | 27.0 | 33.6 | 20.5 | 6.56 | 15.6 | 15.6 | 36.8 | 39.5 | 31.6 | 15.8 | 2.63 | 13.2 | 1.39 |
MAmmoTH | 39.6 | 37.8 | 43.8 | 19.4 | 3.69 | 10.6 | 2.30 | 43.0 | 42.2 | 40.8 | 16.9 | 4.93 | 9.86 | 1.58 |
Llama-2 | 57.6 | 58.6 | 22.9 | 14.3 | 8.10 | 11.4 | 4.76 | 59.6 | 60.3 | 24.4 | 12.8 | 5.77 | 8.97 | 1.72 |
MATHWELL | 69.5 | 69.1 | 24.7 | 10.3 | 5.38 | 7.62 | 1.35 | 71.1 | 70.6 | 24.6 | 8.56 | 4.81 | 7.49 | 1.87 |
MATHWELL Outperforms Existing Models
Our expert annotators scored the 250 outputs from each model for solvability, accuracy, appropriateness, and topic specificity. As shown in Table 2, MATHWELL performs best in each metric, with a 19.9% higher share of outputs with executable code that MaC, 19.9% higher share that have executable code, and 43.4% higher share that have executable code and MaC than the next best model. Table 2 highlights that existing math models are not as capable at this task as Llama-2 and MATHWELL. Specifically, math models struggle with appropriateness and generating executable code, showing that raw coding capability does not necessarily translate into more outputs containing executable code. Table 2 also suggests that QLoRA finetuning on high-quality data is enough to match or exceed math models’ accuracy on this task. Although MATHWELL performs best in topic specificity, the other models do well in this metric, implying that MATHWELL and LLMs can effectively customize word problems to student interests.
@@
Word problems involving multiplication, division, fractions, and decimals are more complex than those only involving addition and subtraction. In Table 3, we assess whether each model can generate MaC questions when using more complicated operations. Questions may contain more than one operation/topic. Solvable questions may require no math operation if they contain the answer in the question (see Appendix E.4.3), so we also report the share of questions containing no operations. To determine if questions with complex operations are accurate and appropriate, we compare the math operations/topics in solvable questions to those in MaC questions. We also report the average number of distinct operations/topics in MaC questions for each model, which is another way to assess question complexity.
Table 3 shows that MATHWELL is the only model for which the share of MaC questions for each math operation/topic is within two percentage points of that for solvable questions, providing evidence that MATHWELL can generate MaC questions regardless of the complexity of the operation. MATHWELL is also the least likely to generate problems that require no operations and has the highest average total operations, two other pieces of evidence that suggest MATHWELL generates high-quality, complex problems.
In our evaluation, we do not prompt models for specific operations. Under these conditions, MATHWELL generates more problems containing addition and subtraction relative to the other models. In turn, there is a concern that MATHWELL’s performance in Table 2 could be due to it generating simple questions for this experiment, which may be more likely to MaC. To address this concern, we conduct two additional analyses reported in Appendix C: 1) logistic regressions showing MATHWELL’s higher MaC relative to the other models holds when controlling for math operations and 2) a summary of accuracy by operation showing MATHWELL is the only model for which accuracy does not substantially differ by operation and remains above 90% for each operation.
@@ 4.2 Automatic Evaluation
Model | PPL | BF1 | GSM BF1 | MaC AL | NDC |
LLEMMA | 3.82 (0.10) | 84.3 | 84.1 | 50.9 (2.89) | 8.41 (0.09) |
MAmmoTH | 2.76 (0.03) | 86.0 | 84.7 | 44.4 (1.15) | 8.25 (0.08) |
Llama-2 | 2.52 (0.03) | 85.5 | 84.3 | 49.8 (1.19) | 8.20 (0.07) |
MATHWELL | 2.44 (0.03) | 85.5 | 84.2 | 54.1 (0.97) | 8.23 (0.08) |
MATHWELL Outputs Have Low Perplexity
Solvability | Accuracy | Appropriateness | ||||||||||||||||
Annotated Test | 78.9 | 61.9 | 86.7 | 88.0 | 87.4 | 0.701 | 81.6 | 53.8 | 87.2 | 92.2 | 89.6 | 0.649 | 82.3 | 63.1 | 86.0 | 93.5 | 89.6 | 0.693 |
LLEMMA | 48.8 | 49.8 | 48.7 | 90.2 | 63.2 | 0.542 | 59.0 | 47.1 | 62.5 | 89.7 | 73.7 | 0.492 | 43.4 | 50.0 | 41.8 | 90.2 | 57.1 | 0.450 |
MAmmoTH | 84.4 | 52.5 | 87.4 | 95.9 | 91.4 | 0.700 | 93.5 | 57.9 | 95.7 | 97.6 | 96.6 | 0.669 | 69.1 | 54.0 | 69.6 | 96.6 | 80.9 | 0.529 |
Llama-2 | 82.4 | 53.1 | 84.9 | 96.2 | 90.2 | 0.619 | 85.2 | 51.6 | 89.8 | 94.1 | 91.9 | 0.677 | 77.6 | 53.7 | 82.2 | 92.4 | 87.0 | 0.579 |
MATHWELL | 85.2 | 54.3 | 90.1 | 93.7 | 91.9 | 0.599 | 94.6 | 62.7 | 97.7 | 96.8 | 97.2 | 0.790 | 83.0 | 57.8 | 88.6 | 92.2 | 90.4 | 0.711 |
MATHWELL Outputs Are Similar to Human-written Questions
Like Jiao et al. (2023) and Zhou et al. (2023), we use BERTScore Zhang et al. (2020) to compute the semantic similarity of our synthetic data and compare it to existing datasets. A lower BERTScore for synthetic questions relative to existing datasets would imply they are less similar to each other than human-written questions, while a higher score would imply they are more similar. We also use BERTScore to compare our synthetic data to MathInstruct GSM8K to identify if they are similar to human-written questions. We calculate MathInstruct GSM8K’s within-dataset BERTScore as a reference to compare each source against. As shown in Tables 4 and 5, across models, SGSM subsets and MathInstruct GSM8K comparisons, BERTScores are similar, suggesting the questions we generate and the data we release are similar to human quality.
MATHWELL Outputs Longer MaC Questions
Longer token length may signal increased question complexity, as longer word problems often contain more information and mathematical operations than shorter ones. Comparing the average length of all questions to the length of MaC questions can determine if MaC questions are shorter or simpler. As shown in Table 5, MATHWELL’s MaC questions are the longest of the models considered, suggesting its MaC problems may be more complex than those from other models. As shown in Table 9, MATHWELL is also the only model for which MaC length is within a token of the overall average, providing evidence that its MaC questions are no simpler than its average question.
MATHWELL Outputs Have Appropriate Readability
Reading level is one way to automatically assess if math word problems are written appropriately. Like Norberg et al. (2023), we use FKGL and NDC to evaluate reading level. FKGL calculates reading level as a function of the total words, total sentences, and total syllables in a piece of text, with the score representing the U.S. grade level of the text Aggarwal ; Flesch (1948); Kincaid et al. (1975). Negative FKGL scores are possible and denote text that is easy to read due to having short words and sentences. NDC calculates readability as a function of sentence length and the proportion of words not contained in a list of 3,000 common English words Aggarwal ; Chall and Dale (1995).
@@
Figure 3 compares the FKGL distribution of SGSM questions to the three other existing datasets that are appropriate for grade school students. Fewer of SGSM’s questions than other existing datasets are written at a grade level beyond 8th grade. This suggests that the dataset captures the intended age range and complements existing datasets by including questions more appropriate for struggling readers. Figure 4 compares the FKGL distribution of MATHWELL outputs to the other models considered. While MATHWELL has a similar average FKGL, it is the only model that does not generate questions beyond an 8th grade reading level, providing evidence it creates more readable questions for K-8 students.
4.3 Question Classifier Performance
As shown in Table 6, despite weighting the loss, each classifier has a lower balanced accuracy than its overall accuracy. The classifiers perform best on output from models derived from Llama-2, although performance for all models tends to be lower than on the test dataset. Excluding LLEMMA and the appropriateness classifier’s precision on MAmmoTH outputs, each classifier has fairly high precision, recall and F1, suggesting they are effectively able to label solvable, accurate, and appropriate questions. The classifiers tend to have low ROC AUC across sources, suggesting they do not perform well at different prediction cutoffs. As a whole, Table 6 provides evidence that text classifiers can learn to label MATHWELL outputs, but future research should use more balanced datasets to improve their ability to label unsolvable, inaccurate, and inappropriate questions.
@@ 5 Conclusions
We explore context-free word problem generation and create MATHWELL, which generates a question/answer pair based only on an optional topic. To train our model, we generate synthetic data and use expert annotators to identify a high-quality training subset. We release SGSM, a synthetic dataset of 20,490 question/answer pairs for use in future research. Our evaluations show that MATHWELL outperforms other open-source LLMs at context-free word problem generation and that SGSM is of comparable quality to existing math QA data. These findings suggest that context-free word problem generation is a feasible and practical alternative to traditional context-dependent generators. Future research should train context-free word problem generators that can create questions aligned with specific math topics and grade levels.
@@
One important limitation of MATHWELL is that it is not designed to generate questions aligned with pre-specified grade levels and mathematical operations/topics, which we chose to leave to future research due to the high cost of annotating questions for these characteristics but would improve the model’s utility in classroom settings. Additionally, MATHWELL is trained and evaluated for generating word problems/solutions for K-8 students only; therefore, we do not recommend using it to generate question/answer pairs for other grade levels or for other tasks.
@@ Acknowledgements
Appendix A Important Criteria for Context-free Word Problem Generator Training Data
The characteristics we find most important for training context-free word problem generators are PoT solutions written as Python functions and questions that are written at an appropriate mathematical difficulty for K-8 students. Regarding the former, when we modified our prompt to ask for a Python function solution instead of a Python code solution, the share of question/answer pairs with executable code from an early version of MATHWELL increased from 18.9% to 29.0%. For appropriateness, when we used the GSM-Hard Gao et al. (2023) dataset as part of MATHWELL’s training data, the model often generated questions with large numbers that are inappropriate for K-8 students. As shown in Table 1, the SGSM Train subset is the only math QA dataset that has these two characteristics.
Appendix B Annotated Data Characteristics
Our annotated dataset consists of 4,234 question/answer pairs with human annotations for solvability, accuracy, appropriateness, and MaC. The data are comprised of the 3,234 word problem dataset used to generate training data for MATHWELL in addition to the 250 evaluation set for each model described in Section 4. Based on our annotations, 81.1% of the question/answer pairs are solvable, 86.5% have accurate solutions, 67.3% are appropriate, and 57.4% meet all criteria.
Appendix C Additional Analyses
C.1 Logistic Regression for Predicting MaC
As shown in Table 7, the coefficients for all models for MaC remain negative and statistically significant relative to MATHWELL when controlling for the type of mathematical operation. This finding supports the assertion that MATHWELL is more capable of generating MaC questions regardless of the operation considered, even if it is less likely to generate questions from the more complex mathematical operations.

### v4

@@ 4 Experiments
| Solvable Questions | MaC Questions | ||||||||||||||
| Model | Add. | Sub. | Mult. | Div. | Frac. | Dec. | No Ops | Add. | Sub. | Mult. | Div. | Frac. | Dec. | Total Ops | |
| API | GPT-4 Turbo | 53.2 | 44.7 | 61.6 | 25.3 | 0.80 | 4.64 | 1.27 | 54.3 | 45.2 | 59.9 | 25.4 | 1.02 | 3.05 | 1.89 |
| GPT-3.5 Turbo | 35.3 | 28.5 | 44.3 | 36.2 | 3.17 | 23.1 | 1.81 | 36.3 | 29.9 | 39.5 | 35.7 | 3.18 | 21.0 | 1.66 | |
| Public | LLEMMA | 34.4 | 27.0 | 33.6 | 20.5 | 6.56 | 15.6 | 15.6 | 36.8 | 39.5 | 31.6 | 15.8 | 2.63 | 13.2 | 1.39 |
| MAmmoTH | 39.6 | 37.8 | 43.8 | 19.4 | 3.69 | 10.6 | 2.30 | 43.0 | 42.2 | 40.8 | 16.9 | 4.93 | 9.86 | 1.58 | |
| Llama-2 | 57.6 | 58.6 | 22.9 | 14.3 | 8.10 | 11.4 | 4.76 | 59.6 | 60.3 | 24.4 | 12.8 | 5.77 | 8.97 | 1.72 | |
| MATHWELL (Ours) | 69.5 | 69.1 | 24.7 | 10.3 | 5.38 | 7.62 | 1.35 | 71.1 | 70.6 | 24.6 | 8.56 | 4.81 | 7.49 | 1.87 | |
We evaluate MATHWELL using both human evaluations with domain experts and automatic evaluations. We compare 250 randomly-selected MATHWELL outputs to those of both open and closed source existing models. For closed-source models, we evaluate GPT-3.5 Turbo and GPT-4 Turbo. GPT-3.5 represents a closed-source model with similar capability to Llama-2 (70B) Touvron et al. (2023), while GPT-4 is a much more capable model OpenAI et al. (2024). For open-source models, we evaluate LLEMMA (34B) Azerbayev et al. (2023), MAmmoTH (70B) Yue et al. (2023), and Llama-2 (70B) Touvron et al. (2023), which are recent and popular. LLEMMA and MAmmoTH are two different approaches to training math LLMs, as MAmmoTH is trained as a QA model Yue et al. (2023), while LLEMMA is trained more generally using a corpus of math content Azerbayev et al. (2023). Llama-2 serves as a baseline to ensure task-specific finetuning improves performance. Each model is prompted using examples from SGSM and asked to create a question/answer pair on a random topic K-8 students are interested in (see Appendix G for a full list of topics used during prompting). Our 250 evaluation sample for each model is roughly 2.5-5 times larger than the human evaluation samples used in other word problem generator studies Jiao et al. (2023); Koncel-Kedziorski et al. (2016); Niyarepola et al. (2022); Qin et al. (2024); Zhou and Huang (2019); Zhou et al. (2023); Zong and Krishnamachari (2023).
@@ 4.1 Human Evaluation
MATHWELL Matches SOTA in Human Evaluation Criteria
Expert annotators scored the 250 outputs from each model for solvability, accuracy, appropriateness, and topic specificity. As shown in Table 3, MATHWELL performs best among open-source models in each metric, with a 19.9% higher share of outputs with executable code that MaC, 19.9% higher share that have executable code, and 43.4% higher share that have executable code and MaC than the next best open-source model. MATHWELL also significantly outperforms GPT-3.5 in human evaluation criteria and matches 94.9% of the performance of GPT-4 in MaC, even outperforming the model in accuracy and appropriateness and nearly closing the gap between open and closed source models on these criteria.
Table 3 highlights that existing open-source math models are not as capable at this task as Llama-2 and MATHWELL. Specifically, the math models struggle with appropriateness and generating executable code, showing that raw coding ability does not necessarily translate into more outputs containing executable code. Table 3 also suggests that QLoRA finetuning on high-quality data is enough to match SOTA models’ performance on human evaluation criteria for this task. Although MATHWELL performs best in topic specificity, the other models do well in this metric, implying that MATHWELL and LLMs can effectively customize word problems to student interests. Lastly, Table 3 highlights that open-source models have a large gap in the share of question/answer pairs that contain executable code relative to GPT-3.5, the best performing model in this metric (see Appendix I for discussion on GPT-4’s performance on this metric).
@@
Word problems involving multiplication, division, fractions, and decimals are more complex than those only involving addition and subtraction. In Table 3, we assess whether each model can generate MaC questions when using more complicated operations. Questions may contain more than one operation/topic. Solvable questions may require no math operation if they contain the answer in the question (see Appendix J.6.3), so we also report the share of questions containing no operations. To determine if questions with complex operations are accurate and appropriate, we compare the math operations/topics in solvable questions to those in MaC questions. We also report the average number of distinct operations/topics in MaC questions for each model, which is another way to assess question complexity.
Table 3 shows that GPT-4 and MATHWELL are the only models for which the share of MaC questions for each math operation/topic is within two percentage points of that for solvable questions, showing MATHWELL can generate MaC questions regardless of the complexity of the operation. GPT-4 and MATHWELL are also the least likely to generate problems that require no operations and have the highest average total operations, two other pieces of evidence that suggest MATHWELL generates high-quality, complex problems.
In our evaluation, we do not prompt models for specific operations. Under these conditions, MATHWELL generates more problems containing addition and subtraction relative to the other models. In turn, there is a concern that MATHWELL’s performance in Table 3 could be due to it generating simple questions for this experiment, which may be more likely to MaC. To address this concern, we conduct two additional analyses reported in Appendix D: 1) logistic regressions showing MATHWELL’s higher MaC relative to the other models except GPT-4 remains statistically significant when controlling for math operations and 2) a summary of accuracy by operation showing GPT-4 and MATHWELL are the only models for which accuracy does not substantially differ by operation and remains above 90% for each operation.
@@ 4.2 Automatic Evaluation
MATHWELL Outputs Have Low Perplexity
MATHWELL Outputs Are Similar to Human-written Questions
Like Jiao et al. (2023) and Zhou et al. (2023), we use BERTScore Zhang et al. (2020) to compute the semantic similarity of our synthetic data and compare it to existing datasets. A lower BERTScore for synthetic questions relative to existing datasets would imply they are less similar to each other than human-written questions, while a higher score would imply they are more similar. We also use BERTScore to compare our synthetic data to MathInstruct GSM8K to identify if they are similar to human-written questions. We calculate MathInstruct GSM8K’s within-dataset BERTScore as a reference to compare each source against. As shown in Tables 4 and 5, across models, SGSM subsets and comparisons, BERTScores are similar, suggesting the data we release are similar to human quality.
@@
| Model | PPL | BF1 | GSM BF1 | MaC AL | NDC | |
| API | GPT-4 Turbo | 2.50 (0.03) | 85.4 | 84.6 | 66.8 (2.62) | 8.26 (0.07) |
| GPT-3.5 Turbo | 2.64 (0.03) | 85.6 | 84.7 | 49.9 (1.16) | 8.58 (0.07) | |
| Public | LLEMMA | 3.82 (0.10) | 84.3 | 84.1 | 50.9 (2.89) | 8.41 (0.09) |
| MAmmoTH | 2.76 (0.03) | 86.0 | 84.7 | 44.4 (1.15) | 8.25 (0.08) | |
| Llama-2 | 2.52 (0.03) | 85.5 | 84.3 | 49.8 (1.19) | 8.20 (0.07) | |
| MATHWELL (Ours) | 2.44 (0.03) | 85.5 | 84.2 | 54.1 (0.97) | 8.23 (0.08) |
MATHWELL Outputs Longer MaC Questions than Existing Open-source Models
Longer token length may signal increased question complexity, as longer word problems often contain more information and mathematical operations than shorter ones. Comparing the average length of all questions to the length of MaC questions can determine if MaC questions are shorter or simpler. As shown in Table 5, MATHWELL’s MaC questions are the longest of the open-source models considered, suggesting its MaC problems may be more complex than those from other models. As shown in Table 8, MATHWELL and GPT-4 are also the only models for which MaC length is within a token of the overall average, providing evidence that its MaC questions are no simpler than its average question. While GPT-4 outputs longer questions than MATHWELL on average, MATHWELL’s question length is still on par with those of the human-written grade school math datasets reported in Table 1.
MATHWELL Outputs Have More Appropriate Readability
Reading level is an automatic way to assess if math word problems are written appropriately. Like Norberg et al. (2023), we use FKGL and NDC to evaluate reading level. FKGL calculates reading level as a function of the total words, total sentences, and total syllables in a piece of text, with the score representing the U.S. grade level of the text Aggarwal ; Flesch (1948); Kincaid et al. (1975). Negative FKGL scores are possible and denote text that is easy to read due to having short words and sentences. NDC calculates readability as a function of sentence length and the proportion of words not contained in a list of 3,000 common English words Aggarwal ; Chall and Dale (1995).
@@
Figure 3 compares the FKGL distribution of SGSM questions to the three other existing datasets that are appropriate for grade school students. Fewer of SGSM’s questions than other existing datasets are written at a grade level beyond 8th grade. This suggests that the dataset captures the intended age range and complements existing datasets by including questions more appropriate for struggling readers. Figure 4 compares the FKGL distribution of MATHWELL outputs to the other models considered. As shown in the figure, MATHWELL has the lowest FKGL and is the only model that does not generate questions beyond an 8th grade reading level, providing evidence it creates more readable questions for K-8 students.
@@ 5 Conclusions
We explore age-appropriate, educational context-free word problem generation and generate a high-quality synthetic dataset verified by human experts to train models for this task. We demonstrate the quality of this dataset by using it to further finetune MATHWELL, the first context-free math word problem generator that generates a question/answer pair based only on an optional topic. Our evaluations show that MATHWELL outperforms other open-source LLMs at context-free word problem generation, nearly matching the performance of GPT-4 on this task while generating questions with a more appropriate reading level. Further, we show that our dataset, SGSM, is of comparable quality to existing math QA data. These findings suggest that context-free word problem generation is a feasible and practical alternative to traditional context-dependent generators. Future research should train context-free word problem generators that can create questions aligned with specific math topics and grade levels and further develop automated classification methods to reduce human annotation costs.
@@
One important limitation of MATHWELL is that it is not designed to generate questions aligned with pre-specified grade levels and mathematical operations/topics, which we chose to leave to future research due to the high cost of annotating questions for these characteristics but would improve the model’s utility in classroom settings. The high cost of human evaluation to ensure age-appropriate, educational outputs is another limitation of automatic word problem generation broadly. In Appendix F, we explore training text classifiers to automatically score MATHWELL outputs. We show existing models can learn some features important for automatic classification, but need refinement in order to correctly classify unsolvable, inaccurate, or inappropriate questions. We hope these results and our large annotated dataset motivate future research in automatic classification efforts. Additionally, MATHWELL is trained and evaluated for generating word problems/solutions for K-8 students only; therefore, we do not recommend using it to generate question/answer pairs for other grade levels or for other tasks.
@@ Acknowledgements
Appendix A Finetuning Details
A.1 Initial Finetuning
We re-format MathInstruct GSM8K Yue et al. (2023) into an Alpaca-style Taori et al. (2023) instruction dataset of question/answer pairs with a standard prompt asking the model to generate a grade school math word problem. To the best of our knowledge, MathInstruct GSM8K is the only dataset with Python code solutions and English questions that are all appropriate for K-8 students Yue et al. (2023). We also include MathInstruct MATH and TheoremQA PoT Yue et al. (2023) question/answer pairs with a standard prompt asking the model to generate a challenge math problem to expose the model to more complex code and further promote mathematical reasoning. We then finetune Llama-2 (70B) on this dataset of 25,926 rows using QLoRA Dettmers et al. (2023) for 4,250 steps. We use two A100 GPUs, a learning rate of 1e-4, LoRA modules at every model layer, per-device batch size of 1, and a 3% warm-up ratio.
A.2 Secondary Finetuning on Expert-annotated Data
Inspired by recent works that iteratively finetune LLMs Guo et al. (2024); Wang et al. (2024), including OpenAI’s FeedME approach Zheng et al. (2023), we next aim to evaluate and improve how well MATHWELL meets our three main criteria: solvability, accuracy, and appropriateness. However, these properties are unannotated in prior works. Therefore, we few-shot prompt the finetuned model and generate 3,234 problems. We then acquire annotations for solvability, accuracy, and appropriateness from volunteers with previous K-12 teaching experience/training. Further details about the annotation process and annotator agreement are in Appendix L. These annotations identified 1,906 problems that are simultaneously solvable, accurate, and appropriate, or meet all criteria (MaC). We then finetune MATHWELL on these MaC outputs for 1,250 more steps.
Appendix B Important Criteria for Context-free Word Problem Generator Training Data
In addition to high-quality grammar and problems that are similar to those students encounter in the classroom, the characteristics we find most important for training context-free word problem generators are PoT solutions written as Python functions and questions that are written at an appropriate mathematical difficulty for K-8 students. Regarding the former, when we modified our prompt to ask for a Python function solution instead of a Python code solution, the share of question/answer pairs with executable code from an early version of MATHWELL increased from 18.9% to 29.0%. For appropriateness, when we used the GSM-Hard Gao et al. (2023) dataset as part of MATHWELL’s training data, the model often generated questions with large numbers that are inappropriate for K-8 students. As shown in Table 1, the SGSM subset is the only math QA dataset that has these two characteristics.
Appendix C Annotated Data Characteristics
Our annotated dataset consists of 4,734 question/answer pairs with human annotations for solvability, accuracy, appropriateness, and MaC. The data are comprised of the 3,234 word problem dataset used to generate training data for MATHWELL in addition to the 250 evaluation set for each model described in Section 4. Based on our annotations, 82.2% of the question/answer pairs are solvable, 87.2% have accurate solutions, 68.6% are appropriate, and 58.8% meet all criteria.
Appendix D Additional Analyses
D.1 Logistic Regression for Predicting MaC
As shown in Table 6, the coefficients for all models except GPT-4 for MaC remain negative and statistically significant relative to MATHWELL when controlling for the type of mathematical operation. This finding supports the assertion that MATHWELL is more capable than these models of generating MaC questions regardless of the operation considered, even if it is less likely to generate questions from the more complex mathematical operations. Table 6 also shows that the difference in MaC performance between MATHWELL and GPT-4 remains statistically insignificant even when controlling for the impact of operations, highlighting that both models perform similarly on this task.